# CogCoM: Train Large Vision-Language Models Diving into Details through Chain of Manipulations

### v2

@@ 4.2 Training
This stage further trains the model to align with human preferences on solving practical visual problems. We fuse the produced CoM data with 3 types of corpus, including MultiInstruct (Xu et al., 2022), LLaVAR (Zhang et al., 2023b), and ShareGPT4V (Chen et al., 2023c), referring the abilities of instruction-following, texts-recognizing, and detailed-captioning. This fusion results in a total of 570K $(I,Q,A)$ samples, where the answer $A$ in CoM data consists of multiple turns. For the training data of CoM, we randomly prepend a lunching prompt [^4] See Appendix D.1 for examples. $P^{\mathcal{M}}$ to questions $Q=P^{\mathcal{M}}+Q$ asking models to optionally use manipulations for the adaption of explicitly eliciting. We empirically show that the model can effectively learn the evidential visual reasoning by ingesting this portion of CoM data. We train model with 14,000 iterations and a batch size of 160, where the learning rate reaches $10^{-5}$ after 280 steps of warm-up and then decays linearly. The parameters of 6.5B visual experts are trained with the objective of next token prediction. These two stages of training result in our standard version of CogCoM involving both chat and reasoning capabilities. More training details are available at Appendix D.2.
@@ 5 Experiment
To quantitatively validate the suitability and efficiency of the proposed method, we conduct experiments on 9 benchmarks corresponding to 4 categories of multimodal capabilities, as well as on a newly constructed testbed that includes the evidential reasoning paths with a keypoints-aware metric. Following previous works, we train two generalist versions of CogCoM for adapting to the different scenarios of Visual Question Answering and Visual Grounding, and evaluate the standard version with a qualitative analysis (Hwang et al., 2023). We also evaluate the time complexity.
@@
  • Visual Grounding. Visual grounding evaluates the crucial abilities of VLMs on meticulous position understanding. We evaluate our model on 3 standard benchmarks, RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016).
  • General Multimodal Capabilities & Hallucination. We also evaluate on a general multimodal benchmark, MM-Vet (Yu et al., 2023b), and a hallucination detection benchmark POPE (Li et al., 2023c), to investigate the helpfulness of visual reasoning.
@@ 5.1 Experiments on Detailed VQA
VLMs have demonstrated the well-known superiority in visual scenes with salient content understanding. We evaluate the effectiveness of CogCoM on VQAs on detailed understanding, which typically require models to perform multiple actions (find, read) or multiple reasoning steps (recognizing and then calculating). Following previous studies (Wang et al., 2023b), we train our model obtained from the first-phase of stage-1 on a mixture of data, including an instruction corpus of MultiInstruct, 13 publicly available VQA datasets (only using training set), a newly created VQA dataset built through promoting GPT-4V (OpenAI, 2023b) for image-oriented question-answer generation, and the automatically generated 70K CoM corpus. This training results in a generalist VQA model incorporating CoM reasoning. For all existing VQA tasks, we directly prompt CogCoM with given questions and examine the correctness of outputted answers.
@@
Table 1: Performance on Visual Question Answering benchmarks, where the results labeled with refer to the few-shot setting. CogCoM achieves SOTA across the board, and demonstrates the effectiveness on the visual reasoning and scene texts recognition benchmarks.
@@ 5.1.1 GQA, TextVQA, ST-VQA, TallyVQA
GQA is a compositional VQA benchmark with diverse reasoning questions coming from semantic functional programs. TallyVQA is an objects counting benchmark with human-annotated complex counting questions involving challenging non-zero counterparts. TextVQA and ST-VQA are two texts understanding benchmarks requiring models to answer questions through textual cues on images. We use the official evaluation scripts for GQA and TallyVQA, which calculate the accuracy score by the Exact Matching (EM) between model predictions and answers. For TextVQA and ST-VQA, we submit our model predictions to the official online websites for calculating the accuracy with VQA Score metric (Antol et al., 2015).
@@ Results
As the results shown in Table 2, CogCoM achieves the state-of-the-art performance in comparison with all generalist models, and achieves significant improvements over the baseline model. Specifically, compared to the baseline model, our model achieves up to 5.97 and 9.0 percentage points improvement on the benchmarks that requires complex reasoning and detailed recognition, respectively. On GQA and TextVQA, CogCoM also obtains comparable results with the large-scale specialist SOTAs. This result demonstrates the effectiveness of the proposed approach in solving details recognition problem.
@@ 5.1.2 Experiments for Reasoning Accuracy and Time Complexity
Due to the lack of resource, we build CoM-test, a benchmark with evidential reasoning chains on the TextVQA test set based on the proposed data generation pipeline, and also introduce a keypoints-aware metric to validate the correctness of reasoning paths (see Appendix C.3 for detailed statistics). We also evaluate the time complexity for model generation on a held-out benchmark, MM-Vet.
Refer to captionRefer to caption
Figure 6: Left: Results on a reasoning testbed CoM-test shows CogCoM achieves satisfactory performance with only 70K training data and 2K steps. Right: Results on MM-Vet shows that CogCoM produces comprehensive reasoning content without incurring excessive time overhead.
@@ 5.2 Experiments on Visual Grounding
Shikra-7B (Chen et al., 2023b) 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
Shikra-13B (Chen et al., 2023b) 87.83 91.11 81.81 82.89 87.79 74.41 82.64 83.16
@@ Results
As shown in Figure 2, CogCoM achieves the best performance in 6 out of all 8 sub-sets. Based on the training with a mixture of broad capabilities, this result indicates that our model exhibits a superior grounding abilities while offers potential to solve a variety of tasks.
@@ 5.3 Experiments on General Multimodal Evaluation and Hallucination Examination
LLaVA-1.5 (Liu et al., 2023a) Vicuna-13B 36.3 84.5
CogVLM (Wang et al., 2023b) Vicuna-7B 45.5 87.2
@@ 6 Conclusion
This paper studies the problems presented by the conclusive alignment training of VLMs, and proposes a mechanism, Chain of Manipulations (CoM), that enables VLMs to solve problems step-by-step by actively manipulating visual inputs as evidence. We realize this methodology by proposing (1) a flexible data structure, (2) an efficient data generation framework capable of producing abundant samples, (3) a memory-based architecture compatible with existing VLMs, and (4) a training process for versatile capabilities. We also annotate 6K graphical math samples with reasoning chains to facilitate the advancement of VLMs in solving mathematical problems. Experiments on 9 public benchmarks show that our trained 17B general VLM can produce informative reasoning content while achieving superior performance on diverse multimodal problems.
Appendix A Related Works
A.1 Large Vision-Langauge Models as Foundations
Most of LVLMs rely on the training on publicly available image-caption pairs, including ALIGN (Jia et al., 2021), MSCOCO (Lin et al., 2014), VG Krishna et al. (2017), CC3M Sharma et al. (2018), CC12M (Changpinyo et al., 2021), SBU (Ordonez et al., 2011), LAION2B (Schuhmann et al., 2022), LAION400M Schuhmann et al. (2021). Starting from Flamingo (Alayrac et al., 2022), a series of LVLMs have focused on training the adaptation layers to align the visual representation to the frozen LLMs on a mixture of image-text pairs with the above corpus, including BLIP2 Li et al. (2023b), KOSMOS Huang et al. (2023b), and OpenFlamingo (Awadalla et al., 2023). Inspired by success of instruction tuning in LLMs (Wang et al., 2022c), a line of works have devoted efforts to build vision-oriented instruction-answer pairs through GPT4 and train models for imitation, such as LLAVA (Liu et al., 2023b), Otter (Li et al., 2023a), VisionLLM (Wang et al., 2023a), MultiInstruct (Xu et al., 2022), Lynx (Zeng et al., 2023), InstructBLIP (Dai et al., ), CleverFlamingo (Chen et al., 2023a) and StableLLaVA (Li et al., 2023d). Recently, researchers have proven the efficiency of developing LVLMs with two stages of training, the first stage of abundant pretraining on image-caption pairs and the second stage of alignment on image-question-answer triples, such as PALI (Chen et al., 2022), PaLI-X (Chen et al., 2023d), Qwen-VL (Bai et al., 2023), and CogVLM Wang et al. (2023b).
A.2 Large Vision-Language Models with Reasoning
To further enhance the ability of LVLMs in solving high-level visual problems, research focusing on various aspects of reasoning is attracting broad attention. We simply divide existing studies into tree broad categories. The first line of research focus on enhance train models with a mastery of cross-modal grounded reasoning, where grounded instruction-following supervision is build through public visual grounding dataset or GPT4-V for training, including KOSMOS-2 (Peng et al., 2023), Shikra (Chen et al., 2023b), and GPT4ROI (Zhang et al., 2023a). The second aspect of efforts have been devoted into promoting models to understand artificial visual scenes, such as figures, charts, and receipts. These studies includes CogAgent (Hong et al., 2023) and CHARTVE (Huang et al., 2023a). Some other studies address the crucial problem of hallucination in LVLMs with counterfactual or interpretable reasoning (Yu et al., 2023a; Yin et al., 2023). V* (Wu & Xie, 2023) also contributes efforts to enhance the details recognition of VLMs based the LLM-guided searching process.
Appendix B Limitation and Impact
@@
Though we try to develop an accurate and robust framework that engages remarkable LLM to provide basic solving steps, adopts reliable visual tools to obtain visual contents, and then acquires feasible paths based on traversal, there are still limitations in our methodology that we hope to improve in the future. First, We find that the diversity of linguistic solving steps is insufficient, and the inaccuracy of visual tools (e.g., the rough granularity of grounding boxes, OCR failures on slant letters) will lead to a large amount of negative paths (effectively utilizing these paths would beneficial). We suggest to promote these limitations with dedicate prompts and improved visual tools. Second, our current model re-input the manipulated images with a set of hard prompts, which may bring speed losses. This is expected to be improved by implementing the physical manipuations into the calculations in vector space. This work presents a general visual reasoning mechanism that alleviate the problems caused by existing conclusion-alignment training for VLMs, introduces a data production framework involving LLMs and visual tools as reliable annotators, and devises a memory-based compatible VLM architecture. We expect this work to bring three benefits to the community. First, the proposed visual reasoning mechanism may push the progress of VLMs in solving complex visual problems. Second, the introduced data production framework may be applied to widespread training scenarios to promote the development of current data-driven machine learning. Third, we hope that the memory-based architecture will be helpful for VLMs in multi-turn long contexts.
Appendix C Details of Data Production

### v3

@@ 4.2 Training
This stage further trains the model to align with human preferences on solving practical visual problems. We fuse the produced CoM data with 3 types of corpus, including MultiInstruct (Xu et al., 2022), LLaVAR (Zhang et al., 2023b), and ShareGPT4V (Chen et al., 2023b), referring the abilities of instruction-following, texts-recognizing, and detailed-captioning. This fusion results in a total of 570K $(I,Q,A)$ samples, where the answer $A$ in CoM data consists of multiple turns. For the training data of CoM, we randomly prepend a lunching prompt [^4] See Appendix F.1 for examples. $P^{\mathcal{M}}$ to questions $Q=P^{\mathcal{M}}+Q$ asking models to optionally use manipulations for the adaption of explicitly eliciting. We empirically show that the model can effectively learn the evidential visual reasoning by ingesting this portion of CoM data. We train model with 14,000 iterations and a batch size of 160, where the learning rate reaches $10^{-5}$ after 280 steps of warm-up and then decays linearly. The parameters of 6.5B visual experts are trained with the objective of next token prediction. These two stages of training result in our standard version of CogCoM involving both chat and reasoning capabilities. More training details are available at Appendix F.2.
@@ 5 Experiment
To quantitatively validate the suitability and efficiency of the proposed method, we conduct experiments on 9 benchmarks corresponding to 4 categories of multimodal capabilities, as well as on a newly constructed testbed that includes the CoM reasoning paths with a keypoints-aware metric. Following previous works, we train two generalist versions of CogCoM to adapt to different scenarios: Visual Question Answering and Visual Grounding, and we evaluate the standard version through a qualitative analysis (Hwang et al., 2023). We also evaluate the time complexity of our model.
@@
  • Visual Grounding. Visual grounding evaluates the crucial abilities of VLMs in meticulous position understanding. We evaluate our model on 3 standard benchmarks, RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016).
  • General Multimodal Capabilities & Hallucination. We also evaluate on a general multimodal benchmark, MM-Vet (Yu et al., 2023b), and a hallucination detection benchmark POPE (Li et al., 2023d), to investigate the helpfulness of visual reasoning.
@@ 5.1 Experiments on Detailed VQA
VLMs have demonstrated well-known superiority in visual scenes with salient content understanding. We evaluate the effectiveness of CogCoM on VQAs for detailed understanding, which typically require models to perform multiple actions (find, read) or multiple reasoning steps (recognizing and then calculating). Following previous studies (Wang et al., 2023a), we train our model obtained from the first-phase of stage-1 on a mixture of data, including an instruction corpus of MultiInstruct, 13 publicly available VQA datasets (only using training set), a newly created VQA dataset built through promoting GPT-4V (OpenAI, 2023b) for image-oriented question-answer generation, and the automatically generated 70K CoM corpus. This training results in a generalist VQA model incorporating CoM reasoning. For all existing VQA tasks, we directly prompt CogCoM with given questions and examine the correctness of outputted answers.
@@
Table 1: Performance on VQA benchmarks, where the results with refer to the few-shot setting.
@@ 5.1.1 GQA, TextVQA, ST-VQA, TallyVQA
GQA is a compositional VQA benchmark with diverse reasoning questions coming from semantic functional programs. TallyVQA is an object counting benchmark with human-annotated complex counting questions involving challenging non-zero counterparts. TextVQA and ST-VQA are two texts understanding benchmarks requiring models to answer questions through textual cues on images. We use the official evaluation scripts for GQA and TallyVQA, which calculate the accuracy score by the Exact Matching (EM) between model predictions and answers. For TextVQA and ST-VQA, we submit our model predictions to the official online websites for calculating the accuracy with VQA Score metric (Antol et al., 2015).
@@ Results
As the results shown in Table 2, CogCoM achieves SOTA performance compared to all generalist models, and achieves significant improvements over the baseline model. Specifically, compared to the baseline model, our model achieves up to 5.97 and 9.0 percentage points improvement on the benchmarks that require complex reasoning and detailed recognition, respectively. On GQA and TextVQA, CogCoM also obtains comparable results with the large-scale specialist SOTAs. This result demonstrates the effectiveness of the proposed approach in solving details recognition problem.
@@ 5.1.2 Experiments for Reasoning Accuracy and Time Complexity
Due to the lack of resources, we build CoM-test, a benchmark with CoM reasoning chains on the TextVQA test set based on the proposed data generation pipeline, and also introduce a keypoints-aware metric to validate the correctness of reasoning paths (see Appendix E.3 for detailed statistics). We also evaluate the time complexity for model generation on a held-out benchmark, MM-Vet.
Figure 6: Left: Results on a reasoning testbed CoM-test show CogCoM achieves satisfactory performance with only 70K training data and 2K steps. Right: Results on MM-Vet show that CogCoM produces comprehensive reasoning content without incurring excessive time overhead.
@@ 5.2 Experiments on Visual Grounding
Shikra-7B (Chen et al., 2023a) 87.01 90.61 80.24 81.60 87.36 72.12 82.27 82.19
Shikra-13B (Chen et al., 2023a) 87.83 91.11 81.81 82.89 87.79 74.41 82.64 83.16
@@ Results
As shown in Figure 2, CogCoM achieves the best performance in 6 out of all 8 sub-sets. Based on the training with a mixture of broad capabilities, this result indicates that our model exhibits a superior grounding ability while offers potential to solve a variety of tasks. In addition, CogCoM achieves performance on par with the specialist SOTAs and surpasses the ONE-PEACE with a leading performance on the test-A from RefCOCO+. This result demonstrates that under a generalized training integrating multiple capabilities, our model engages grounding as a foundational skill and cultivate the capability to accomplish complex problems.
@@ 5.3 Experiments on General Multimodal Evaluation and Hallucination Examination
Monkey (Li et al., 2024) Qwen-7B 36.2 -
LLaVA-1.5 (Liu et al., 2023a) Vicuna-13B 39.4 85.0
CogVLM (Wang et al., 2023a) Vicuna-7B 45.5 87.2
@@ Results
5.4 Ablation Experiment for Training w/wo CoM Data
We conduct experiments on our generalist VQA version model CogCoM-chat, where we removed the 70K CoM training data from the training procedure and keep all other settings unchanged (i.e., using the corpus of public VQAs and 500K training data from MultiInstruct, LLaVAR, and ShareGPT4V). The results on three typical benchmarks are shown in Table4. We can see that our model benefits significantly from the produced CoM training data on these benchmarks that require detailed recognition, multimodal reasoning and mathematical capabilities to achieve substantial improvements.
      Model       TextVQA       MMVet       MathVista
      CogCoM (wo)       64.5       45.9       34.8
      CogCoM (w)       71.1 ($\uparrow$ 6.6)       46.1 ($\uparrow$ 0.2)       35.7 ($\uparrow$ 0.9)
Table 4: Ablation experimental results for training CogCoM with/without 70K CoM training data.
6 Related Works
Most of LVLMs rely on the training on publicly available image-caption pairs, including ALIGN (Jia et al., 2021), MSCOCO (Lin et al., 2014), VG Krishna et al. (2017), CC3M Sharma et al. (2018), CC12M (Changpinyo et al., 2021), SBU (Ordonez et al., 2011), LAION2B (Schuhmann et al., 2022), LAION400M Schuhmann et al. (2021). Starting from Flamingo (Alayrac et al., 2022), a series of LVLMs have focused on training the adaptation layers to align the visual representation to the frozen LLMs on a mixture of image-text pairs with the above corpus, including BLIP2 Li et al. (2023b), KOSMOS Huang et al. (2023b), and OpenFlamingo (Awadalla et al., 2023). Inspired by success of instruction tuning in LLMs (Wang et al., 2022c), a line of works have devoted efforts to build vision-oriented instruction-answer pairs through GPT4 and train models for imitation, such as LLAVA (Liu et al., 2023b), Otter (Li et al., 2023a), VisionLLM (Wang et al., 2023b), MultiInstruct (Xu et al., 2022), Lynx (Zeng et al., 2023), InstructBLIP (Dai et al., ), and StableLLaVA (Li et al., 2023c). Recently, researchers have proven the efficiency of developing LVLMs with two stages of training, the first stage of abundant pretraining on image-caption pairs and the second stage of alignment on image-question-answer triples, such as PALI (Chen et al., ), PaLI-X (Chen et al., 2023c), Qwen-VL (Bai et al., 2023), and CogVLM Wang et al. (2023a).
To further enhance the ability of LVLMs in solving high-level visual problems, research focusing on various aspects of reasoning is attracting broad attention. We simply divide existing studies into three broad categories. The first line of research focuses on enhance train models with a mastery of cross-modal grounded reasoning, where grounded instruction-following supervision is build through public visual grounding dataset or GPT4-V for training, including KOSMOS-2 (Peng et al., 2023), Shikra (Chen et al., 2023a), and GPT4ROI (Zhang et al., 2023a). The second aspect of efforts have been devoted to promoting models to understand artificial visual scenes, such as figures, charts, and receipts. These studies includes CogAgent (Hong et al., 2023) and CHARTVE (Huang et al., 2023a). Some other studies address the crucial problem of hallucination in LVLMs with counterfactual or interpretable reasoning (Yu et al., 2023a; Yin et al., 2023). V* (Wu & Xie, 2023) also contributes efforts to enhance the details recognition of VLMs based on the LLM-guided searching process.
@@ 7 Conclusion
This paper studies the problems presented by the conclusive alignment training of VLMs, and proposes a mechanism, Chain of Manipulations (CoM), that enables VLMs to solve problems step-by-step by actively manipulating visual inputs as evidence. We realize this methodology by proposing (1) a flexible data structure, (2) an efficient data generation framework capable of producing abundant samples, (3) a memory-based architecture compatible with existing VLMs, and (4) a training process for versatile capabilities. We also annotate 7K graphical math samples with reasoning chains to facilitate the advancement of VLMs in solving mathematical problems. Experiments on 9 public benchmarks show that our trained 17B general VLM can produce informative reasoning content while achieving superior performance on diverse multimodal problems.
Acknowledgments
This work is supported by National Natural Science Foundation of China (No. 62277033), and National Natural Science Foundation of China (No. 62476150). Thanks the support from National Engineering Laboratory for Cyberlearning and Intelligent Technology, Beijing Key Lab of Networked Multimedia, and Zhipu AI. This work is also supported by the Natural Science Foundation of China (NSFC) for Distinguished Young Scholar 62425601, and the Natural Science Foundation of China (NSFC) 62495063.
Appendix A Discussion with Closely Related Works
The efforts of ViperGPT (Surís et al., 2023) share the same basic idea with our work that decomposes complex visual problems into reasoning steps. In comparison with their training-free framework combining external VLMs using a code LLM, we focus on training an end-to-end VLM to enhance its visual reasoning ability to solve complex problems. V* (Wu & Xie, 2023) is a concurrent work who also solves problems by progressively acquiring visual cues. Their two-part framework first utilizes a VQALLM model to list visual objects, followed by a dedicated search model to acquire the objects. On the other hand, our approach focuses on using one model to actively perform reasoning and to identify or mark the most useful visual information, which may offer the potential for solving more complex reasoning tasks in addition to the detailed identification, such as the challenging geometric math problems. DualFocus (Cao et al., 2024) was released around the same time as ours. They also construct a training dataset that includes intermediate cues (bounding boxes) and trained the model with two stages. Compared to their work, our CoM training places more emphasis on answering questions in a single reasoning process and marking images to assist in solving complex problems. VPD (Hu et al., 2024) converts programs obtained from LLM and execution engine into CoT and distills them into the VLM, enabling the model to reason when solving visual problems. This approach is similar to our method on visual reasoning. However, our ultimate goal is to train VLMs to solve complex visual problems by actively reasoning (i.e., CoT) and manipulating (e.g., zooming in or drawing auxiliary lines) images which is consistent with the human behavior in realistic scenarios. VisProg (Gupta & Kembhavi, 2023) relies on in-context learning of LLMs to generate programs, which are then executed to get final answers. Ferret-v2 (Zhang et al., 2024) add the Dense Referring and Dense Detection tasks into the training stages and adopts an additional DINO encoder to enhance the fundamental visual capability of VLMs. The study from (Bhattacharyya et al., ) trains a VLM to combine the low-level visual features with high-level inferences to reason and generate the final response on videos, where the low-level features are introduced by surrogate tasks during training.
Appendix B Discussion for Future Works
Humans solve difficult questions with a period of thinking before answering. In contrast, LLMs/VLMs generate outputs immediately after prompting tokens. The Chain of Thought (CoT) mechanism serves as an effective substitute for this thinking process. Similar to human reasoning, a backtracking mechanism could provide a reliable approach for forming correct and concise reasoning paths. However, as the integration of backtracking results in a high time complexity, developing an efficient backtracking strategy that improves the accuracy of reasoning paths without sacrificing performance is a crucial direction for future work.
Appendix C Limitation and Impact
@@
Though we build a robust framework with remarkable LLMs and reliable visual tools, there are still limitations. First, the diversity of solving steps is still insufficient, and inaccuracies in visual tools (e.g., coarse grounding boxes, OCR errors on slanted text) result in many negative paths, which could be better utilized. We propose improving these issues with dedicated prompts and enhanced visual tools. Second, re-inputting manipulated images with fixed prompts can slow down the process. This can be optimized by incorporating physical manipulations directly into the vector space calculations. This work presents a general visual reasoning mechanism that alleviates the problems caused by existing conclusion-alignment training for VLMs, introduces a data production framework involving LLMs and visual tools as reliable annotators, and devises a memory-based compatible VLM architecture. We expect this work to bring three benefits to the community. First, the proposed visual reasoning mechanism may push the progress of VLMs in solving complex visual problems. Second, the introduced data production framework may be applied to widespread training scenarios to promote the development of current data-driven machine learning. Third, we hope that the memory-based architecture will be helpful for VLMs in multi-turn long contexts.
Appendix D Additional Analysis
D.1 The Number of Inference Turns and the Quality Control
We have detailed the statistics for the training and testing data in the Appendix E.3, including the total number of reasoning chains, the average number of reasoning steps, and the average number of manipulation types. To clarify the average and maximum numbers of the multi-image inference turns, we have also conducted statistics on the number of turns divided by multiple images. The results are as follows: in the training data source from TextVQA and ST-VQA which may involve generating new images such as through zooming, the average number of turns is 1.42, and the maximum number of turns is 7 (we restrict the maximum number of turns to 4 during training to prevent OOM). In the test set of TextVQA, our model produced an average of 1.54 turns involving multiple images. It is worth noting that not every image requires manipulations such as zooming, and some can be answered through reasoning with evidence or direct observation.
During the collection of our 70K CoM training data, we discard the wrongly recognized data (i.e., we refer to these data as the negative paths in our paper), as these data can not terminate to the golden answer node during the DFS traversal. We used this filtering strategy to ensure that only the correct data capable of reaching the golden answer (i.e., positive paths) was included in the 70K training data. Since we recursively search for the answer by following the intermediate path composed of grounding boxes until we reach the leaf answer node, this approach generally produces the correct path, except in a few cases where the grounding boxes may be too large. Most of the data we have constructed relies on grouping results as intermediate reasoning steps. After our manual verification, we found that most of the reasoning paths are indeed correct. Therefore, most of the generated CoM training samples can be guaranteed to be error-free. As using Reinforcement Learning to penalize the negative paths during training is another optimization strategy, we look forward to utilizing these negative paths as negative rewards in future work.
D.2 The Distribution of the Successful and Unsuccessful paths
Since most of the questions in the constructed dataset ask about the details of images, we have compiled the distribution of image classes for the reasoning data sourced from TextVQA (which includes an image classes label). The statistics are now presented in Figure 7 of the revised paper. Specifically, we analyze the distribution of image classes where a successful reasoning path was not obtained using GPT-4 and visual tools, and among those classes, the distribution of image classes where a successful reasoning path was achieved. To facilitate display, we show the top 200 image classes. To improve visualization, we applied $\log_{10}(\cdot)$ to the positive frequency values on the y-axis, excluding zero. We find that: (1) For most image classes, the proportions of successful and unsuccessful paths are approximately consistent, indicating that the image category does not have a significant impact in this problem. (2) The success rate for constructing reasoning paths is higher for common and prominent objects, such as “person” and “vehicle”, while the success rate is lower for smaller, less common objects like ”necklace” and “kettle.” This suggests that the lower success rate for such objects is due to the limitations in the accuracy of the tools used to find the positive path.
Figure 7: Distributions based on top200 image classes from unsuccessful paths, where the positive frequencies on yaxis are taken as $\log_{10}(\cdot)$ to facilitate the visualization.
Appendix E Details of Data Production